OpenStudy (anonymous):

What is the y intercept of the line that is parallel to 2x+3y=4 and contains the point (6,-2)?

OpenStudy (anonymous):

parallel line has slope \(-\frac{2}{3}\) use the point slope formula to find the equation of your line and the y-intercept

OpenStudy (anonymous):

\[y+2=-\frac{2}{3}(x-6)\] is a start, formula is \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

2x+3y=4 3y=-2x+4 y=(-2/3)x+4 your line's slope is -2/3, so the perpendicular line is gonna have slope 3/2 (negative reciprocal) so you have y=(3/2)x+b for your second line, so plug in the point -2=(3/2)6+b -2=9+b b=-11 your y intercept is -11.

OpenStudy (anonymous):

thats not one of the choices

OpenStudy (anonymous):

you sure you typed the problem in correct? you're looking for a y intercept right, not an x? and the point is (6,-2) and the equation is right?

OpenStudy (anonymous):

yea i did

OpenStudy (anonymous):

it can be -2 1 2 3

OpenStudy (anonymous):

ohhh sh*t. sorry. you said parallel, not perpendicular. i looked at the first guy's answer and we both read your problem wrong. so the first equation has slope -2/3, so your second equation will have slope -2/3 too. y=(-2/3)x+b -2=(-2/3)6+b -2=-4+b b=2 2 is your answer, sorry for that.

OpenStudy (anonymous):

i am pretty sure that is what i said

OpenStudy (anonymous):

\[y+2=-\frac{2}{3}(x-6)\] \[y+2=-\frac{2}{3}x+4\] \[y=-\frac{2}{3}x+2\]